OpenStudy (ttop0816):

i want to see if my answer is right! please help! If the ordinate is three more than twice the abscissa and x {-1, 0, -}, which of the following sets of ordered pairs satisfies the given conditions? {(-1, 1), (0, 3), (-1/3, 7/3)}<-- i chose this! {(-1, 1), (0, 3), (-1/3, 11/3)} {(-1, -1), (0, 3), (-1/3, 1)}

Directrix (directrix):

The y-term is the ordinate; the x-term is the abscissa. So, y = 3 + 2x. I checked out the points in the option you indicated as correct and agree with you. By "checked out," I mean I calculations such as: y = 3 + 2x 1 = ? 2*(-1) + 3 @ttop0816